# SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM

### 02 Mar 2024 v3

@@ 4.3 Evaluation of Semantic Segmentation
table 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches. It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes. In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%. Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based approaches. Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation. In contrast, NeRF-based methods often struggle to recognize individual objects and typically require complex muti-level model designs and extensive feature fusion. Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
Table 3: Quantitative comparison of our method against existing semantic NeRF-based SLAM methods on the Replica dataset Straub et al. (2019). The baselines are limited to four scenes as their results are reported only for these. For each scene, we compute the average mIoU score by comparing the rendered and the ground-truth 2D semantic image in the training view. Our method significantly outperforms the NeRF-based approaches, achieving mIoU scores over 90%.
Methods Avg. mIoU [%$\uparrow$ Room0 [%$\uparrow$ Room1 [%$\uparrow$ Room2 [%$\uparrow$ Office0 [%$\uparrow$
NIDS-SLAM 82.37 82.45 84.08 76.99 85.94
[…]
@@ 4.4 Evaluation of Key Frame Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset Dai et al. (2017a). The rendered views display the reconstructed bike and guitar captured within the trajectory. Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable. Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the accuracy of a previously well-established scene. A case in point is scene0000 from the ScanNet dataset Dai et al. (2017a), where objects such as bike and guitar are revisited at early and late stages in the trajectory. Keyframes from later in the sequence, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions of these objects. fig. 3 illustrates the novel-view evaluation for scene0000. In comparison to SplaTAM, depicted in the upper row, our method delivers more accurate reconstruction outcomes. The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved. Our method facilitates the selection of keyframes based on geometric and semantic constraints, incorporating uncertainty weighting during the optimization of selected keyframes. This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
@@ 4.5 Scene Manipulation
Figure 4: The case study on scene manipulation in room0 of the Replica dataset Straub et al. (2019). We show the capabilities for object removal and transformation by specifying semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a jar and flowers, as well as moving and rotating them.
@@ 4.6 Ablation Study
We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al. (2017a) to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
@@ 4.6.1 Effect of Multi-channel Optimization
table 4 shows the ablation study on multi-channel parameter optimization. The results reveal that our optimization strategy can significantly improve the localization and mapping performance. Specifically, the system without appearance color cannot provide rendered views, whereas camera pose and depth can still be estimated by leveraging depth and semantic input. The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision. Furthermore, the absence of an input semantic map disables 3D semantic segmentation and remarkably diminishes the performance of tracking and mapping. Additionally, the silhouette threshold, essential for assessing scene visibility, is crucial for the system stability. Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where appearance, depth, and semantic supervision are removed.
@@
without color image 7.44 24.59 68.19
without depth map 47.66 40.47 15.14 54.52
without semantic map 9.15 13.81 17.52
without silhouette threshold 29.12 357.48 12.06 28.07
with multi-channel inputs 6.66 11.26 19.47 70.27
@@ 4.6.2 Effect of Keyframe Optimization
table 5 presents the results of keyframe selection ablation. Our two-level keyframe selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance. Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
@@
without geometric threshold 6.66 15.55 19.21 68.93
without semantic threshold 8.44 12.89 17.84 69.85
without uncertainty weighting 6.87 11.43 18.72 70.12
@@ 5.1 Limitations
Appendix A Experiment Settings
In this section, we outline the experimental setup and hyperparameters applied in our studies. The experiments were conducted on NVIDIA A100-40GB GPUs.
A.0.1 SGS-SLAM
By default, both mapping and tracking operations are conducted for each frame. During the tracking phase, we set the silhouette visibility threshold, $T_{\rm sil}$, to 0.99. The multi-channel optimization involves three parameters: $\lambda_{D}=1.0$ for depth, $\lambda_{C}=0.5$ for colors, and $\lambda_{S}=0.05$ for semantic loss, with the semantic loss weight being comparatively low due to the typical noisiness of real-world semantic labels. Throughout the tracking, the multi-channel Gaussian parameters remain constant, adjusting only the camera parameters with a learning rate of 2e-3 for transition. Key-frames are initially chosen at intervals of every 5 frames, then refined based on geometric and semantic criteria. The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7. The maximum number of keyframes per frame is limited to 25, considering the computation speed. The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series. In mapping process, the silhouette threshold $T_{\rm sil}$ is adjusted to 0.5. The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$. Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and log scale at 1e-3. Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
The mapping and tracking iteration steps are specific to each dataset, In the case of the Replica dataset Straub et al. [2019], the number of iterations for tracking and mapping are set to 40 and 60. For the ScanNet dataset Dai et al. [2017a], tracking and mapping are set to 120 and 40. In the enhanced ScanNet++ dataset Yeshwanth et al. [2023], where the camera transition is large between each frame, the tracking and mapping iterations are adjusted to 220 and 50.
A.0.2 Baselines
We adhere to the default configurations for each baseline as reported in their papers. The evaluation metrics for tracking and mapping are consistent with those applied to our method. For baselines whose implementations are not publicly available, we present the results as reported in their papers.
Appendix B Additional Experiment Results
We provide additional quantitative analysis of camera tracking in section B.1. The visualization of semantic segmentation compared with NeRF-based method is presented in section B.2. More qualitative novel view rendering results are illustrated in section B.3. We compared our method with Vox-Fusion Yang et al. [2022], NICE-SLAM Zhu et al. [2022], Co-SLAM Wang et al. [2023], ESLAM Johari et al. [2023], and Point-SLAM Sandström et al. [2023] for ATE RMSE evaluation. For 3D semantic segmentation, we visualized the comparison with DNS-SLAM Li et al. [2023a].
B.1 Camera Tracking
In this section, we break down the quantitative analysis on ATE RMSE [cm] on Replica Straub et al. [2019], ScanNet Dai et al. [2017a], and ScanNet++ Yeshwanth et al. [2023] datasets. table 6, table 7, and table 8 present the evaluation our SGS-SLAM against baseline models on each dataset. Our method of estimating camera poses by directly optimizing the gradient on dense photometric loss achieves state-of-the-art tracking performance on datasets with high-quality RGB-D images. In particular, on the ScanNet++ dataset Yeshwanth et al. [2023], where there is a large camera transition between successive frames, NeRF-based methods like ESLAM failed to track. Conversely, SGS-SLAM demonstrated robust tracking performance.
Table 6: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for each scene of the Replica dataset Straub et al. [2019].
Methods Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
Vox-Fusion 3.09 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94
[…]
Table 7: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for the selected scenes on the ScanNet dataset Dai et al. [2017a].
Methods Avg. $0$ 0059 0106 0169 0181 0207
Vox-Fusion 26.90 68.84 24.18 8.41 27.28 23.30 9.41
[…]
Table 8: Quantitative comparison of ATE RMSE [cm] between our method and the baseline for the selected scenes on the ScanNet++ dataset Yeshwanth et al. [2023].
Methods Avg. [cm]$\downarrow$ 8b5caf3398 [cm]$\downarrow$ b20a261fdf [cm]$\downarrow$
ESLAM 170.06 185.15 156.96
[…]
B.2 Semantic Segmentation
In this section, the outcomes of semantic segmentation on the Replica dataset Straub et al. [2019] are visualized and compared with DNS-SLAM Li et al. [2023a], a NeRF-based approach. As illustrated, our method offered accurate and detailed segmentation, whereas DNS-SLAM faces challenges in edges due to the over-smoothing issue of NeRF.
Figure 5: Qualitative comparison of our method and DNS-SLAM Li et al. [2023a] for semantic segmentation from the Replica dataset Straub et al. [2019]. The visualization outcomes of DNS-SLAM Li et al. [2023a] are obtained from its paper. The frames of the training view are chosen based on the results presented in DNS-SLAM. Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
B.3 Novel View Rendering
We present additional results of novel view rendering using our method across the Replica Straub et al. [2019], ScanNet Dai et al. [2017a], and ScanNet++ Yeshwanth et al. [2023] datasets, with comparisons to ESLAM Johari et al. [2023]. Visualizations are provided in fig. 6, fig. 7, fig. 8, and fig. 9 with semantic segmentation outcomes. Our method consistently delivers high-quality rendering results for both synthesized and real-world datasets. Notably, on the challenging real-world ScanNet++ dataset, ESLAM Johari et al. [2023] struggled to reconstruct the scene. By contrast, SGS-SLAM provides accurate high-fidelity scene reconstructions along with precise segmentation outcomes. Note that the ground-truth segmentation labels are retrieved from the ground-truth mesh at the instance level, and therefore, our results also show instance-level segmentation.
Figure 6: The visualization of novel view rendering between the ESLAM Johari et al. [2023] and our method on the Replica dataset Straub et al. [2019]. The ground-truth novel views are captured from meshes provided by the dataset.
Figure 7: The visualization of novel view rendering between the baseline and our method using the ScanNet dataset Dai et al. [2017a]. The ground-truth novel views are captured from meshes. SGS-SLAM exhibits rendering of high fidelity and outperforms the NeRF-based ESLAM Johari et al. [2023]. In contrast to the ground-truth mesh, our method demonstrates robust mapping in areas where the ground-truth mesh presents holes.
Figure 8: The visualization of 3D semantic segmentation results of SGS-SLAM, as applied to the novel views selected in fig. 7. Note that the rendering results exhibit minor variations in scene objects due to the use of a modified semantic dataset from ScantNet. For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black). Furthermore, we introduce extra labels, like guitar, bag, and basket, to enhance the quality of scene reconstruction.
Figure 9: The visualization of novel view rendering between the baseline and our method using the ScanNet++ dataset Yeshwanth et al. [2023]. The ground-truth novel views are captured from meshes. SGS-SLAM demonstrates superior rendering quality, while ESLAM Johari et al. [2023] suffers from significant tracking errors and fails to reconstruct the map. In addition, our method also offers accurate instance-level segmentation outcomes.
B.4 Scene Manipulation
In this section, we visualize scene manipulation results by grouping the Gaussians using the semantic mask. As shown in fig. 10, for object removal, we can directly erase the Gaussians associated with the editing target, such as removing the table while preserving all the items on it. In addition, we can group objects by selecting their semantic masks and applying translation and rotation, such as moving and rotating both the table and the above objects to a different place.
Figure 10: The visualization of scene manipulation by grouping Gaussians via semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.

### 13 Mar 2024 v4

@@ 4.3 Evaluation of Semantic Segmentation
Tab. 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches. It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes. In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%. Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based approaches. Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation. In contrast, NeRF-based methods often struggle to recognize individual objects and typically require complex muti-level model designs and extensive feature fusion. Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
Table 3: Quantitative comparison of our method against existing semantic NeRF-based SLAM methods on the Replica dataset [34]. The baselines are limited to four scenes as their results are reported only for these. For each scene, we compute the average mIoU score by comparing the rendered and the ground-truth 2D semantic image in the training view. Our method significantly outperforms the NeRF-based approaches, achieving SOTA mIoU scores over 90%.
  Methods Avg. mIoUnormal-$\uparrow$
Room0 [%$\uparrow$
Room1 [%$\uparrow$
Room2 [%$\uparrow$
Office0 [%$\uparrow$
NIDS-SLAM
82.37
82.45
84.08
76.99
85.94
DNS-SLAM
84.77
88.32
84.90
81.20
84.66
SNI-SLAM
87.41
88.42
87.43
86.16
87.63
Ours 92.72 92.95 92.91 92.10 92.90
@@ 4.4 Evaluation of Keyframe Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset [4]. The rendered views display the reconstructed objects such as bike, fridge, garbage bin, and guitar from novel views. Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable. Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the previously well-established scenes. A case in point is scene0000 from the ScanNet dataset [4], where objects such as bike and guitar are revisited at early and late stages in the trajectory. Keyframes from later sequences, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions. Fig. 3 illustrates the novel view evaluation for scene0000. In comparison to ESLAM [15] and SplaTAM [16], which are based on NeRF and 3D Gaussians, our method delivers more accurate reconstruction outcomes. The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved. Our method facilitates the selection of keyframes based on geometric and semantic constraints, incorporating uncertainty weighting during the optimization of selected keyframes. This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable keyframse with high uncertainty to significantly altering the earlier accurately reconstructed map.
@@ 4.5 Scene Manipulation
Figure 4: The case study on scene manipulation in room0 of the Replica dataset [34]. We show the capabilities for object removal and transformation by specifying semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a jar and flowers, as well as moving and rotating them.
@@ 4.6 Ablation Study
We perform the ablation of SGS-SLAM on the scene0000 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
@@ 4.6.1 Effect of Multi-channel Optimization
Tab. 4 shows the ablation study on multi-channel parameter optimization. The results reveal that our optimization strategy can significantly improve the localization and mapping performance. Specifically, the system without appearance color cannot provide rendered views, whereas camera pose and depth can still be estimated by leveraging depth and semantic input. The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision. Furthermore, the absence of an input semantic map disables 3D semantic segmentation and remarkably diminishes the performance of tracking and mapping. Additionally, the silhouette threshold, essential for assessing scene visibility, is crucial for the system stability. Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where appearance, depth, and semantic supervision are removed.  means the metric is inapplicable.
@@
without color image ($C_{\rm pix}$) 7.44 24.59 68.19
without depth map ($D_{\rm pix}$) 47.66 40.47 15.14 54.52
without semantic map ($S_{\rm pix}$) 9.15 13.81 17.52
without silhouette threshold ($Sil_{\rm pix}$) 29.12 357.48 12.06 28.07
with multi-channel optimization 6.18 11.26 19.47 70.27
@@ 4.6.2 Effect of Keyframe Optimization
Tab. 5 presents the results of keyframe selection ablation. Our two-level keyframe selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance. Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
@@
without geometric threshold ($T_{\rm geo}$) 6.66 15.55 19.21 68.93
without semantic threshold ($T_{\rm sem}$) 8.44 12.89 17.84 69.85
without uncertainty weighting ($\mathcal{U}$) 6.87 11.43 18.72 70.12